A scale on a map says that 1 inch represents 60 miles. If the distance between two points on the map is 2 1/2 inches, how many miles are between the two points?
Answer by TimothyLamb(4379)

m/r = 1 in / 60 miles
1/60 = 2.5/r
r = 2.5(60/1)
---
answer:
150 miles
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php